The boxplot of the exam scores for the students in an introductory statistics course is given below. The distribution of theA Ph.D. graduate has applied for a job with two universities: A and B. The graduate feels that she has a 60% chance of receivThe following is an example of a Frequency 1997 2000 1998 1999 Year Stem and leaf plot Pie chart Bar chart Line chart
Answer #1

#1) box plot

I If the longer part is to the left (or below) the median, the data is skewed left. hence above box plot is negatively skewed

# option b is correct

2)

Given:

here offer from university A =P(A)=0.6

and from university B =P(B)=0.5

probabilty of receiving offer from A, given received offer from B=P(A|B)=0.8

hence probability that both universities will make her an offer =P(A &B)

P(A &B)=P(A|B)*P(B)

   =0.8*0.5

=0.4

#hence probability that both universities will make her an offer =0.4

#option c is correct

Q3)

fromm above chart it is example of bar chart

#Bar Chart

option c is correct